A body of A moves an initial velocity of 2.5 m/s and collides with an identical body of which is at rest. Find the final velocity of A and B after collision if this collision is perfectly elastic
1
Expert's answer
2020-11-10T06:52:26-0500

Let's consider the momentum conservation law:


"mv_1 = mv_1' + mv_2'"

where "m" is the mass of the bodies, "v_1 = 2.5m\/s" is the velocity of the first body before the collistion and "v_1', v_2'" are velocities of the first and second bodies respectively. Factoring out "m", obtain equation (1):


"v_1 = v_1' + v_2'"

Since the collision is perfectly elastic, let's consider the energy conservation law and obtain equation (2):


"\dfrac{mv_1^2}{2} = \dfrac{mv_1'^2}{2} +\dfrac{mv_2'^2}{2} \\
v_1^2 = v_1'^2 + v_2'^2"

Expressing "v_1'" from the equation (1) and substituting it into the equation (2), obtain:


"v_1' = v_1-v_2'\\
v_1^2 = ( v_1-v_2')^2 + v_2'^2\\
v_1^2 = v_1^2 - 2v_1v_2' + v_2'^2 + v_2'^2\\
2v_2'^2 - 2v_1v_2' = 0\\
v_2'(v_2' - v_1) = 0"

There are two possible solutions:


"v_2' = 0, \space v_2' = v_1"

The first one doesn't fit, for it means that no collision occured. Thus, the answer is:


"v_2' = v_1 = 2.5m\/s"

The velocity of the first object after the collision is then:


"v_1' = v_1 - v_2' = 0m\/s"

Answer. 0 m/s, 2.5 m/s.